OpenStudy (anonymous):

What is the y-intercept of: 1: slope=2, passes through (-5,-17) 2: slope=-5, passes through (2,-20) 3: slope=1/2, passes through (-2,-1)

OpenStudy (anonymous):

y intercept;when x=0

OpenStudy (anonymous):

just plug them each into y=mx+b

OpenStudy (anonymous):

and solve for b in each

OpenStudy (anonymous):

y-y1=m(x-x1) y-(-17)=2(x-(-5) y+17=2x+10 y=2x+10-17 x=0 at the y intercept y=-7